Foxgirl0918:

A cell phone plan has a monthly cost that is shown in the table below. What is the correct statement regarding the average rate of change during the 40-minute time of talk?

shonyce15:

Monthly cost of cell phone for 0 minutes is $19.95 and for 40 minutes of talk is $21.95. It means that average rate of change ( increase of monthly cost for each minute of talk time ) is: ARC = ( 21.95 - 19.95 ) / ( 40 - 0 ) = 2.00 / 40 = $0.05. Answer: C )

Foxgirl0918:

k thank you!!!